Find the equation of the line that passes through the point (4,4) and is parallel to the line y=-x+3
Found 2 solutions by Big Poop, robertb:
Answer by Big Poop(157)
Answer by robertb(5830)

The parallel line has to have the form y = -x + c, where c is an undetermined constant. To find the value of c, use the coordinates of the point.
==> 4 = -4 + c ==> c = 8
==> The equation of the line is y = -x + 8. (Equations for one line are fundamentally the same, so it's up to you to just go from one form to another.)